Dave and Sandra pool their individual savings of R2 000 each per month and deposit this into a money market account earning 5% interest per annum, compounded annually. What will the balance of the account be at the end of five years?


1
Expert's answer
2021-03-16T08:56:24-0400

"principal=4000*12=48000"

"Amount=p(1+R\/n)"nt

P-rincipal

R-interest rate.

n-number of times interest applied per time period.

nt-number of time periods elapsed.

"48000(1+0.05)"5

"=R61261.515"

4000=is the principal paid by two monthly.

48000=is the principal fo in a whole year.

61264.515=is the total amount they will have after five years.( Interest and principal)